### v1

# Birkhoff–Kellogg type results in product spaces and their application to differential systems

Abstract. We provide a new version of the well-known Birkhoff-Kellogg invariant-direction Theorem in product spaces. Our results concern operator systems and give the existence of component-wise eigenvalues, instead of scalar eigenvalues as in the classical case, that have corresponding eigenvectors with all components nontrivial and localized by their norm. We also show that, when applied to nonlinear eigenvalue problems for differential equations, this localization property of the eigenvectors provides, in turn, qualitative properties of the solutions. This is illustrated in two context of systems of PDEs and ODEs. We illustrate the applicability of our theoretical results with two explicit examples.
2020 MSC: Primary 47H10, secondary 34B08, 35J57, 45G15, 47H11.
Key words and phrases. Birkhoff–Kellogg type result, nonlinear eigenvalue problem, operator system, nontrivial solution.
1. Introduction
A celebrated result in Nonlinear Analysis is the Birkhoff-Kellogg invariant-direction Theorem [5]. This theorem has been object of extensive research in the past and finds significant applications in the study of nonlinear eigenvalue problems in infinite-dimensional normed linear spaces, see for example the books [2, 12, 18], the recent papers [7, 15], and references therein. In the version by Krasnosel’skiĭ and Ladyženskiĭ [19], a similar result is set in cones of real Banach spaces, yielding the existence of a pair, constituted by a positive eigenvalue and an eigenvector, the latter localized inside a cone. We stress that a notable, common feature of the two above mentioned theorems is that they provide a localization of the eigenfunction; this localization, in turn, provides qualitative properties of the solution in the context of applications to nonlinear eigenvalue problems for differential equations. In the framework of systems the situation is somewhat more delicate. In fact, a direct application of one of the previous two results in product spaces would provide the existence of an eigenvalue with a corresponding (vectorial) eigenfunction that may have trivial components; this issue has been been highlighted for example in [15, Definition 1].
In this note we present a new version of the Birkhoff–Kellogg Theorem in product spaces, which, instead of a scalar eigenvalue, provides the existence of component-wise eigenvalues that have corresponding eigenvectors with all components nontrivial and localized by their norm. In the context of systems of integral equations and their applications, component-wise eigenvalues have been investigated in Chapter 3 of the book [1], where the authors sought constant-sign eigenvectors by means of topological tools such as the Schauder and the Krasnosel’skiĭ-Guo fixed point theorems, and in the book [14] where the authors sought, also via topological fixed point theory, the existence of positive eigenvectors.
Namely, we study systems of type
$\left\{\begin{array}[]{l}x=\lambda_{1}\,T_{1}(x,y),\\ y=\lambda_{2}\,T_{2}(x,y),\end{array}\right$. (1.1)
where $T=(T_{1},T_{2})$ is a suitable compact operator acting on the Cartesian product of two sets $C_{1}$ and $C_{2}$, which can be as follows:
  1. (1)
    $C_{1}$ and $C_{2}$ are cones;
  2. (2)
    $C_{1}$ is a cone and $C_{2}=X_{2}$ is a infinite dimensional normed space;
  3. (3)
    both $C_{1}=X_{1}$ and $C_{2}=X_{2}$ are infinite dimensional normed spaces.
In the context of fixed point theory, a component-wise approach has been utilized in the past, see for example [3, 4, 16, 17, 23, 24]; here we develop a somewhat analogue theory in the framework of nonlinear spectral theory. In particular, we give quite natural conditions yielding the existence of component-wise eigenvalues $\lambda_{1},\lambda_{2}>0$ and corresponding eigenfunctions $x_{0},y_{0}$, both of prescribed non-zero norm. We also provide further results in the settings (2) and (3), which yield the additional existence of negative eigenvalues with corresponding eigenfunctions.
Our existence results are motivated by the applications; in particular we show how our theory can be applied to differential systems. More in details, firstly we focus on the systems of PDEs
$\left\{\begin{array}[]{ll}-\Delta u=\lambda_{1}f(x,u,v),&\ \text{ in }\Omega,\\ -\Delta v=\lambda_{2}g(x,u,v),&\ \text{ in }\Omega,\\ u=v=0,&\ \text{ on }\partial\,\Omega,\end{array}\right$. (1.2)
where $\Omega\subset\mathbb{R}^{n}$ denotes the open unit ball in $\mathbb{R}^{n}$ and $f,g$ are suitable continuous functions. Problems of this type are well-studied, both in the classical case of the spectral problem, that is with $\lambda=\lambda_{1}=\lambda_{2}$, see for example [6, 9, 16, 22], and in the case of component-wise eigenvalues with possibly different $\lambda_{1},\lambda_{2}$, see for instance [8, 13, 21] and references therein. In Theorem 3.1 below we provide sufficient conditions for (1.2) that ensure the existence of a pair of positive component-wise eigenvalues with eigenfunctions possessing nontrivial components with localized norms. The result is then illustrated in a specific example.
Our second setting of application is a BVP for the system of ODEs,
$\left\{\begin{array}[]{l}u^{\prime\prime}(t)+\lambda_{1}f(t,u(t),v(t))=0,\ t\in[0,1],\\ v^{\prime\prime}(t)+\lambda_{2}g(t,u(t),v(t))=0,\ t\in[0,1],\\ u(0)=u(1)=0=v^{\prime}(0)=v(1)-\frac{1}{2}v^{\prime}(1),\end{array}\right$. (1.3)
where $f,g$ are suitable continuous functions. The BCs that occur in (1.3) have been investigated for a different set of parameters by Lan[20]. We rewrite the system (1.3) in terms of a system of Hammerstein integral equations, and we work in the Cartesian product of a conical shell times a ball in the space of continuous functions. Also in this case, our approach yields the existence of two distinct pairs of component-wise eigenvalues with nontrivial eigenfunctions; this is illustrated in a toy model as well.
Overall results are new from both the theoretical and the applied point of view and complement the ones in [1, 2, 6, 8, 9, 13, 14, 15, 16, 21, 22, 24].
2. Birkhoff-Kellogg type results
We begin this Section by recalling the classical Birkhoff-Kellogg invariant-direction Theorem [5], cf. [11, Theorem 6.1].
Theorem 2.1.
Let $U$ be a bounded open neighbourhood of 0 in an infinite-dimensional normed linear space $(V,\|\ \|)$, and let $T:\partial U\to V$ be a compact map satisfying $\|T(x)\|\geq\alpha$ for some $\alpha>0$ for every $x$ in $\partial U$. Then there exist $x_{0}\in\partial U$ and $\lambda_{0}\in(0,+\infty)$ such that $x_{0}=\lambda_{0}T(x_{0}).$
In the following version by Krasnosel’skiĭ and Ladyženskiĭ [19], cf. [18, Theorem 5.5], a similar result is set in cones of real Banach spaces; we recall that a cone $K$ of a normed linear space $(X,\|\,\|)$ is a closed set with $K+K\subset K$, $\mu K\subset K$ for all $\mu\geq 0$ and $K\cap(-K)=\{0\}$.
Theorem 2.2.
Let $X$ be a real Banach space, $U\subset X$ be an open bounded set with $0\in U$, $K\subset X$ be a cone, $T:K\cap\overline{U}\longrightarrow K$ be compact and suppose that
$\inf_{x\in K\cap\partial\,U}\left\|Tx\right\|>0.$
Then there exist $x_{0}\in K\cap\partial\,U$ and $\lambda_{0}>0$ such that $x_{0}=\lambda_{0}\,Tx_{0}$.
Before stating our results we fix some notation. Let $(X,\|\,\|)$ be a normed linear space and $K$ a cone in $X$. Given $r>0$, by $B_{r}$ we mean the open ball in $X$ centered at the origin and with radius $r$, while by $\overline{B}_{r}$, $\partial\overline{B}_{r}$ we mean the closed disk and its boundary, respectively. Moreover, we denote by $K_{r}=B_{r}\cap K$, and by $\overline{K}_{r}=\overline{B}_{r}\cap K$, resp. $\partial\overline{K}_{r}=\partial\overline{B}_{r}\cap K$, the closure and boundary of $K_{r}$ relative to $K$.
Observe that $\partial\overline{K}_{r}$ is a retract of $\overline{K}_{r}$. An explicit example of such a retraction can be found in [10, Example 3] defined as
$\rho(z)=r\dfrac{z+(r-\left\|z\right\|)^{2}h}{\left\|z+(r-\left\|z\right\|)^{2}h\right\|},\quad z\in\overline{K}_{r},$
where $h\in K\setminus\{0\}$ is fixed.
With abuse of notation (the whole space $X$ is not a cone) we will still denote $X_{r}=B_{r}$, $\overline{X}_{r}=\overline{B}_{r}$, $\partial\overline{X}_{r}=\partial\overline{B}_{r}$, so that, if $X$ is infinite dimensional, again $\partial\overline{X}_{r}$ is a retract of $\overline{X}_{r}$.
Let $X_{1}$ and $X_{2}$ be normed linear spaces and $C_{1}\subset X_{1}$, $C_{2}\subset X_{2}$ such that for each $i\in\{1,2\}$ either
  1. (a)
    $C_{i}=K_{i}$ is a cone; or
  2. (b)
    $C_{i}=X_{i}$ is an infinite dimensional normed space.
The following result is a version of the Birkhoff–Kellogg Theorem in product spaces.
Theorem 2.3.
Let $r_{1},r_{2}$ be positive constants and suppose that
$T=(T_{1},T_{2}):\overline{C}_{1,r_{1}}\times\overline{C}_{2,r_{2}}\longrightarrow C_{1}\times C_{2}$
is a compact map satisfying that
$\inf_{\left\|x\right\|=r_{1},\ \left\|y\right\|\leq r_{2}}\left\|T_{1}(x,y)\right\|>0\quad\text{and}\quad\inf_{\left\|x\right\|\leq r_{1},\ \left\|y\right\|=r_{2}}\left\|T_{2}(x,y)\right\|>0.$ (2.1)
Then there exist $\lambda_{1},\lambda_{2}>0$ and $(x_{0},y_{0})\in\overline{C}_{1,r_{1}}\times\overline{C}_{2,r_{2}}$ with $\left\|x_{0}\right\|=r_{1}$ and $\left\|y_{0}\right\|=r_{2}$ such that
$\left\{\begin{array}[]{l}x_{0}=\lambda_{1}\,T_{1}(x_{0},y_{0}),\\ y_{0}=\lambda_{2}\,T_{2}(x_{0},y_{0}).\end{array}\right$. (2.2)
Proof.
For each $i=1,2$, let us consider a retraction $\rho_{i}:\overline{C}_{i,r_{i}}\rightarrow\partial\overline{C}_{i,r_{i}}$. Now, define the auxiliary map $N=(N_{1},N_{2}):\overline{C}_{1,r_{1}}\times\overline{C}_{2,r_{2}}\rightarrow\overline{C}_{1,r_{1}}\times\overline{C}_{2,r_{2}}$ as
$N(x,y)=\left(r_{1}\dfrac{T_{1}\left(\rho_{1}(x),y\right)}{\left\|T_{1}\left(\rho_{1}(x),y\right)\right\|},r_{2}\dfrac{T_{2}\left(x,\rho_{2}(y)\right)}{\left\|T_{2}\left(x,\rho_{2}(y)\right)\right\|}\right),$
and observe that, by (2.1), $N$ is well-defined. Since $N$ is a compact map, Schauder fixed point theorem (see e.g. [11, §6, Theorem 3.2]) ensures that $N$ has at least one fixed point $(x_{0},y_{0})\in\overline{C}_{1,r_{1}}\times\overline{C}_{2,r_{2}}$. Observe that $N\left(\overline{C}_{1,r_{1}}\times\overline{C}_{2,r_{2}}\right)\subset\partial\overline{C}_{1,r_{1}}\times\partial\overline{C}_{2,r_{2}}$ and so it follows that $(x_{0},y_{0})\in\partial\overline{C}_{1,r_{1}}\times\partial\overline{C}_{2,r_{2}}$, that is,
$x_{0}=r_{1}\dfrac{T_{1}(x_{0},y_{0})}{\left\|T_{1}(x_{0},y_{0})\right\|},\quad y_{0}=r_{2}\dfrac{T_{2}(x_{0},y_{0})}{\left\|T_{2}(x_{0},y_{0})\right\|}.$
[…]
Remark 2.4.
It should be noted that, under the assumptions of Theorem 2.3, the existence of $\lambda_{0}>0$ and $(x_{0},y_{0})\in\overline{C}_{1,r_{1}}\times\overline{C}_{2,r_{2}}$ with $\left\|x_{0}\right\|=r_{1}$ and $\left\|y_{0}\right\|=r_{2}$ solving the equation
$\left\{\begin{array}[]{l}x_{0}=\lambda_{0}\,T_{1}(x_{0},y_{0}),\\ y_{0}=\lambda_{0}\,T_{2}(x_{0},y_{0})\end{array}\right$. (2.3)
cannot be guaranteed. Indeed, consider as Banach spaces $X=Y=\mathbb{R}$, the cones $K_{1}=K_{2}=[0,+\infty)$, $r_{1}=r_{2}=1$ and the continuous function $T:[0,1]\times[0,1]\rightarrow[0,+\infty)\times[0,+\infty)$ given by
$T(x,y)=(T_{1}(x,y),T_{2}(x,y))=(2x+y,x+3y).$
Note that $\inf_{x=1}\left|T_{1}(x,y)\right|=2>0$ and $\inf_{y=1}\left|T_{2}(x,y)\right|=3>0$, but there is no $\lambda\in(0,+\infty)$ such that $(1,1)=\lambda\,T(1,1)=\lambda\,(3,4)$.
Let us focus now on operators defined in the product of a cone times an infinite dimensional normed space. In this case, an additional solution can be obtained.
Theorem 2.5.
Let $K_{1}$ be a cone in the normed linear space $X_{1}$ and $X_{2}$ be an infinite dimensional normed space. Let $r_{1},r_{2}$ be positive constants and suppose that $T=(T_{1},T_{2}):\overline{K}_{1,r_{1}}\times\overline{B}_{r_{2}}\longrightarrow K_{1}\times X_{2}$ is a compact map satisfying that
$\inf_{\left\|x\right\|=r_{1},\ \left\|y\right\|\leq r_{2}}\left\|T_{1}(x,y)\right\|>0\quad\text{and}\quad\inf_{\left\|x\right\|\leq r_{1},\ \left\|y\right\|=r_{2}}\left\|T_{2}(x,y)\right\|>0.$
Then there exist $\lambda_{1,1},\lambda_{2,1},\lambda_{1,2}>0$, $\lambda_{2,2}<0$ and $(x_{0,j},y_{0,j})\in\overline{K}_{1,r_{1}}\times\overline{B}_{r_{2}}$, $j=1,2$, with $\left\|x_{0,j}\right\|=r_{1}$ and $\left\|y_{0,j}\right\|=r_{2}$ such that
$\left\{\begin{array}[]{l}x_{0,j}=\lambda_{1,j}\,T_{1}(x_{0,j},y_{0,j}),\\ y_{0,j}=\lambda_{2,j}\,T_{2}(x_{0,j},y_{0,j}),\end{array}\right.\qquad(j=1,2).$
Proof.
The first solution is ensured by Theorem 2.3. In order to obtain the second one, just consider the map $\tilde{N}:\overline{K}_{1,r_{1}}\times\overline{B}_{r_{2}}\rightarrow\overline{K}_{1,r_{1}}\times\overline{B}_{r_{2}}$ given by
$\tilde{N}(x,y)=\left(r_{1}\dfrac{T_{1}\left(\rho_{1}(x),y\right)}{\left\|T_{1}\left(\rho_{1}(x),y\right)\right\|},-r_{2}\dfrac{T_{2}\left(x,\rho_{2}(y)\right)}{\left\|T_{2}\left(x,\rho_{2}(y)\right)\right\|}\right).$
As a consequence of Schauder fixed point theorem, $\tilde{N}$ has a fixed point $(x_{0,2},y_{0,2})$ located in $\partial\overline{K}_{1,r_{1}}\times\partial\overline{B}_{r_{2}}$, that is,
$\left\{\begin{array}[]{l}x_{0,2}=\lambda_{1,2}\,T_{1}(x_{0,2},y_{0,2}),\\ y_{0,2}=\lambda_{2,2}\,T_{2}(x_{0,2},y_{0,2}),\end{array}\right$.
[…]
Remark 2.6.
Under the assumptions of Theorem 2.3, if both $C_{1}$ and $C_{2}$ are infinite dimensional normed spaces, then there exist four couples of numbers $\lambda_{1},\lambda_{2}$ and associated points $(x_{0},y_{0})\in\overline{B}_{r_{1}}\times\overline{B}_{r_{2}}$ with $\left\|x_{0}\right\|=r_{1}$ and $\left\|y_{0}\right\|=r_{2}$ such that (2.2) holds. Indeed, it suffices to apply the Schauder theorem to each auxiliary map
$N_{j,k}(x,y)=\left((-1)^{j}r_{1}\dfrac{T_{1}\left(\rho_{1}(x),y\right)}{\left\|T_{1}\left(\rho_{1}(x),y\right)\right\|},(-1)^{k}r_{2}\dfrac{T_{2}\left(x,\rho_{2}(y)\right)}{\left\|T_{2}\left(x,\rho_{2}(y)\right)\right\|}\right),\quad j,k\in\{1,2\}.$
3. Some applications to differential systems
We apply now the above results to some classes of systems of BVPs in the context of ODEs and PDEs.
3.1. Eigenvalues for a systems of elliptic PDEs
We begin by illustrating the applicability of Theorem 2.3 in the context of PDEs. In particular, we discuss the existence of eigenvalues and eigenfunctions of quasilinear elliptic systems subject to homogeneous Dirichlet boundary conditions of the form
$\left\{\begin{array}[]{ll}-\Delta u=\lambda_{1}f(x,u,v),&\ \text{ in }\Omega,\\ -\Delta v=\lambda_{2}g(x,u,v),&\ \text{ in }\Omega,\\ u=v=0,&\ \text{ on }\partial\,\Omega,\end{array}\right$. (3.1)
where $\Omega\subset\mathbb{R}^{n}$ denotes the open unit ball in $\mathbb{R}^{n}$, $f:\overline{\Omega}\times\mathbb{R}_{+}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ and $g:\overline{\Omega}\times\mathbb{R}_{+}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ are continuous functions. It is folklore that the system (3.1) can be rewritten in the integral form
$\left\{\begin{array}[]{l}u(x)=\displaystyle\lambda_{1}\int_{\Omega}k(x,y)\,f(y,u(y),v(y))\,dy,\\[8.5359pt] v(x)=\displaystyle\lambda_{2}\int_{\Omega}k(x,y)\,g(y,u(y),v(y))\,dy,\end{array}\right$. (3.2)
where $k$ is the Green’s function associated to the linear PDE
$\Delta u=h(x)\ \text{ in }\Omega,\quad u=0\ \text{ on }\partial\,\Omega,$
with $h$ a continuous function. We make use of the cone of positive functions in the space $\mathcal{C}(\overline{\Omega})$ of continuous functions endowed with the usual supremum norm $\left\|u\right\|_{\infty}=\max_{x\in\overline{\Omega}}\left|u(x)\right|$. Namely, we let
$P:=\{u\in\mathcal{C}(\overline{\Omega}):u\geq 0\}.$
With these ingredients we can state the following Theorem.
Theorem 3.1.
Let $r_{1},r_{2}$ be positive constants and suppose that exist two continuous functions $\underline{f},\underline{g}:\overline{\Omega}\rightarrow\mathbb{R}_{+}$ such that the following conditions hold:
  1. $a)$
    $f(x,u,v)\geq\underline{f}(x)$ on $\overline{\Omega}\times[0,r_{1}]\times[0,r_{2}]$ and
    $\sup_{x\in\overline{\Omega}}\int_{\Omega}k(x,y)\underline{f}(y)\,dy>0;$
  2. $b)$
    $g(x,u,v)\geq\underline{g}(x)$ on $\overline{\Omega}\times[0,r_{1}]\times[0,r_{2}]$ and
    $\sup_{x\in\overline{\Omega}}\int_{\Omega}k(x,y)\underline{g}(y)\,dy>0.$
Then there exist $\lambda_{1},\lambda_{2}>0$ and $(u_{0},v_{0})\in P\times P$ with $\left\|u_{0}\right\|_{\infty}=r_{1}$ and $\left\|v_{0}\right\|_{\infty}=r_{2}$ that satisfy the system (3.2).
Proof.
Let us consider the Banach spaces $X=Y=\mathcal{C}(\overline{\Omega})$ and the cones $K_{1}=K_{2}=P$. By classical PDE theory, the operator
$T=(T_{1},T_{2}):\overline{P}_{r_{1}}\times\overline{P}_{r_{2}}\longrightarrow P\times P,$
defined by
$\begin{array}[]{l}T_{1}(u,v)(x)=\displaystyle\int_{\Omega}k(x,y)\,f(y,u(y),v(y))\,dy,\\[8.5359pt] T_{2}(u,v)(x)=\displaystyle\int_{\Omega}k(x,y)\,g(y,u(y),v(y))\,dy,\end{array}$ (3.3)
is compact. Now let us consider $(u,v)\in P\times P$ such that $\left\|u\right\|_{\infty}=r_{1},\left\|v\right\|_{\infty}\leq r_{2}$. For every $x\in\overline{\Omega}$ we have
$\|T_{1}(u,v)\|_{\infty}\geq T_{1}(u,v)(x)=\displaystyle\int_{\Omega}k(x,y)\,f(y,u(y),v(y))\,dy\geq\displaystyle\int_{\Omega}k(x,y)\,\underline{f}(y)\,dy.$
Then we get
$\|T_{1}(u,v)\|_{\infty}\geq\sup_{x\in\overline{\Omega}}\int_{\Omega}k(x,y)\,\underline{f}(y)\,dy.$ (3.4)
Note that the RHS of (3.10) does not depend on the particular $(u,v)$ chosen. Therefore we obtain
$\inf_{\left\|u\right\|_{\infty}=r_{1},\ \left\|v\right\|_{\infty}\leq r_{2}}\left\|T_{1}(u,v)\right\|_{\infty}\geq\sup_{x\in\overline{\Omega}}\int_{\Omega}k(x,y)\,\underline{f}(y)\,dy>0.$
[…]
In the following example we show the applicability of Theorem 3.1.
Example 3.2.
Take the open set $\Omega=\{(x,y)\in\mathbb{R}^{2}:x^{2}+y^{2}<1\}$ and consider the system
$\left\{\begin{array}[]{ll}-\Delta u=\lambda_{1}(1+x^{2})e^{u}(2+\cos v),&\text{ in }\Omega,\\[8.5359pt] -\Delta v=\lambda_{2}(1+y^{2})(1+v^{2})(2+\sin u),&\text{ in }\Omega,\\[8.5359pt] u=v=0,&\text{ on }\partial\,\Omega.\end{array}\right$. (3.5)
Now fix $r_{1},r_{2}>0$, then the conditions $a)$ and $b)$ of Theorem 3.1 are satisfied with the choice of $\underline{f}(x,y)=\underline{g}(x,y)\equiv 1$, since a direct calculation gives
$\sup_{(x,y)\in\overline{\Omega}}\int_{\Omega}k((x,y),(w,z))\,d(w,z)=\sup_{(x,y)\in\overline{\Omega}}\dfrac{1}{4}(1-x^{2}-y^{2})=\dfrac{1}{4}.$
Note that $(r_{1},r_{2})$ can be chosen arbitrarily in $(0,+\infty)\times(0,+\infty)$, thus we obtain the existence of infinitely many couples of type $(\lambda_{1},\lambda_{2})$, with $\lambda_{1},\lambda_{2}>0$, and associated couples of nonnegative functions $(u_{0},v_{0})$ with prescribed norm that satisfy the system (3.5).
3.2. Eigenvalues for a system of ODEs
Here we apply Theorem 2.5 to the study of eigenvalues and eigenfunctions for the following class of BVPs for systems of ODEs:
$\left\{\begin{array}[]{l}u^{\prime\prime}(t)+\lambda_{1}f(t,u(t),v(t))=0,\ t\in[0,1],\\ v^{\prime\prime}(t)+\lambda_{2}g(t,u(t),v(t))=0,\ t\in[0,1],\\ u(0)=u(1)=0=v^{\prime}(0)=v(1)-\frac{1}{2}v^{\prime}(1),\end{array}\right$. (3.6)
where $f:[0,1]\times\mathbb{R}_{+}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ and $g:[0,1]\times\mathbb{R}_{+}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ are continuous functions.
Note that the system (3.6) can be rewritten as a system of Hammerstein integral equations, namely
$\left\{\begin{array}[]{r}u(t)=\lambda_{1}\displaystyle\int_{0}^{1}k_{1}(t,s)f(s,u(s),v(s))\,ds,\\[8.5359pt] v(t)=\lambda_{2}\displaystyle\int_{0}^{1}k_{2}(t,s)g(s,u(s),v(s))\,ds,\end{array}\right$. (3.7)
where $k_{1}$ and $k_{2}$ are the corresponding Green’s functions, which are given by
$k_{1}(t,s)=\left\{\begin{array}[]{ll}{(1-t)s},&\quad s\leq t,\\ {t(1-s)},&\quad s>t,\end{array}\right$.
and
$k_{2}(t,s)=\frac{1}{2}\left\{\begin{array}[]{ll}1-2t,&\quad s\leq t,\\ 1-2s,&\quad s>t.\end{array}\right$.
In this case we utilize the space $C[0,1]$, endowed with the usual supremum norm $\|u\|_{\infty}:=\max_{t\in[0,1]}|u(t)|$, and work within the product of the cone of positive functions
$K=\bigl{\{}u\in C[0,1]:u\geq 0,\displaystyle{\min_{t\in[1/4,3/4]}}u(t)\geq\frac{1}{4}\|u\|_{\infty}\bigr{\}}$ (3.8)
with the space itself.
With these ingredients we can state the following Theorem.
Theorem 3.3.
Let $r_{1},r_{2}$ be positive constants and suppose that exist two continuous functions $\underline{f},\underline{g}:[0,1]\rightarrow\mathbb{R}_{+}$ such that the following conditions hold:
  1. $a)$
    $f(t,u,v)\geq\underline{f}(t)$ on $1/4,3/4]\times[r_{1}/4,r_{1}]\times[-r_{2},r_{2}]$ and
    $\sup_{t\in[1/4,3/4]}\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)\underline{f}(s)\,ds>0;$
  2. $b)$
    $g(t,u,v)\geq\underline{g}(t)$ on $0,1]\times[0,r_{1}]\times[-r_{2},r_{2}]$ and
    $\int_{0}^{1}\underline{g}(s)\,ds>0.$
Then there exist $\lambda_{1,1},\lambda_{2,1},\lambda_{1,2}>0$, $\lambda_{2,2}<0$ and $(u_{0,j},v_{0,j})\in\overline{K}_{r_{1}}\times\overline{B}_{r_{2}}$, $j=1,2$, with $\left\|u_{0,j}\right\|=r_{1}$ and $\left\|v_{0,j}\right\|=r_{2}$ that satisfy the system (3.7).
Proof.
Let us consider the Banach spaces $X=Y=C[0,1]$ and the cone $K$ as in (3.8). Note that the operator
$T=(T_{1},T_{2}):\overline{K}_{r_{1}}\times\overline{B}_{r_{2}}\longrightarrow K\times C[0,1],$
defined by
$\begin{array}[]{r}T_{1}(u,v)(t):=\displaystyle\int_{0}^{1}k_{1}(t,s)f(s,u(s),v(s))\,ds,\\[8.5359pt] T_{2}(u,v)(t):=\displaystyle\int_{0}^{1}k_{2}(t,s)g(s,u(s),v(s))\,ds.\end{array}$ (3.9)
is compact.
Firstly, let us take $(u,v)\in K\times C[0,1]$ such that $\left\|u\right\|_{\infty}=r_{1},\left\|v\right\|_{\infty}\leq r_{2}$. Note that for every $t\in[1/4,3/4]$ we have $u(t)\geq r_{1}/4$ and, furthermore, we have
$\|T_{1}(u,v)\|_{\infty}\geq T_{1}(u,v)(t)=\displaystyle\int_{0}^{1}k_{1}(t,s)f(s,u(s),v(s))\,ds\\ \geq\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)f(s,u(s),v(s))\,ds\geq\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)\underline{f}(s)\,ds.$
Then we get
$\|T_{1}(u,v)\|_{\infty}\geq\sup_{t\in[1/4,3/4]}\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)\underline{f}(s)\,ds.$ (3.10)
Note that the RHS of (3.10) does not depend on the particular $(u,v)$ chosen. Therefore we obtain
$\inf_{\left\|u\right\|_{\infty}=r_{1},\ \left\|v\right\|_{\infty}\leq r_{2}}\left\|T_{1}(u,v)\right\|_{\infty}\geq\sup_{t\in[1/4,3/4]}\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)\underline{f}(s)\,ds>0.$
Secondly, let us take $(u,v)\in K\times C[0,1]$ such that $\left\|u\right\|_{\infty}\leq r_{1},\left\|v\right\|_{\infty}=r_{2}$. Note that
[…]
Note that the RHS of (3.11) does not depend on the particular $(u,v)$ chosen. Therefore we obtain
$\inf_{\left\|u\right\|_{\infty}\leq r_{1},\ \left\|v\right\|_{\infty}=r_{2}}\left\|T_{2}(u,v)\right\|_{\infty}\geq\displaystyle\int_{0}^{1}-k_{2}(1,s)\underline{g}(s)\,ds=\frac{1}{2}\displaystyle\int_{0}^{1}\underline{g}(s)\,ds>0.$
[…]
In the following example we show the applicability of Theorem 3.3.
Example 3.4.
Consider the system
$\left\{\begin{array}[]{l}u^{\prime\prime}(t)+\lambda_{1}t(1+u^{2}v^{2})=0,\ t\in[0,1],\\ v^{\prime\prime}(t)+\lambda_{2}te^{uv}=0,\ t\in[0,1],\\ u(0)=u(1)=0=v^{\prime}(0)=v(1)-\frac{1}{2}v^{\prime}(1).\end{array}\right$. (3.12)
Now fix $r_{1},r_{2}>0$, then the choice of $\underline{f}(t)\equiv\frac{1}{4}$ gives, by direct calculation,
$\sup_{t\in[1/4,3/4]}\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)\frac{1}{4}\,ds=\frac{3}{128}>0,$
while choosing $\underline{g}(t)=te^{-r_{1}r_{2}}$ yields
$\int_{0}^{1}te^{-r_{1}r_{2}}\,dt=\frac{e^{-r_{1}r_{2}}}{2}>0.$
Then the conditions $a)$ and $b)$ of Theorem 3.3 are satisfied. Furthermore since the pair $(r_{1},r_{2})$ can be chosen arbitrarily in $(0,+\infty)^{2},$ we obtain the existence of two distinct families of uncountably many pairs: both $(\lambda_{1,1},\lambda_{2,1})$, with $\lambda_{i,1}>0$, $i=1,2$, and $(\lambda_{1,2},\lambda_{2,2})$, with $\lambda_{1,2}>0$ and $\lambda_{2,2}<0$, each of them with the associated eigenfunctions of prescribed norms.
Acknowledgements
A. Calamai and G. Infante are members of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). G. Infante is a member of the UMI Group TAA and the “The Research ITalian network on Approximation (RITA)”. J. Rodríguez–López has been partially supported by the VIS Program of the University of Calabria, by Ministerio de Ciencia y Tecnología (Spain), AEI and Feder, grant PID2020-113275GB-I00, and by Xunta de Galicia, grant ED431C 2023/12. This study was partly funded by: Research project of MIUR (Italian Ministry of Education, University and Research) Prin 2022 “Nonlinear differential problems with applications to real phenomena” (Grant Number: 2022ZXZTN2).